# Reasoning AI Performance Degradation in 6G Networks with Large Language Models

### v1

@@ Abstract
6G, AI Model Reasoning, Large Language Model, Chain-of-Thought, Multi-Access TechnologyI Introduction
As technology continues to evolve, network infrastructure is rapidly advancing towards the next-generation framework, notably 6G. Unlike its predecessors, 6G is expected to feature a deep and pervasive integration of Artificial Intelligence (AI) [1]. The native AI integration has the potential to significantly enhance the functionality and efficiency of 6G networks,, offering powerful capabilities in terms of connectivity, reliability, and intelligent decision-making. However, the performance of AI models within these networks is critical. Any degradation in AI performance can profoundly affect network operations and the services they support. Various factors can contribute to AI performance degradation, such as data quality issues, algorithmic biases, and sudden changes in network conditions. Moreover, the complex and dynamic nature of 6G can deteriorate these challenges, as AI models may struggle to adapt swiftly to the high-speed, voluminous data flows that characterize such advanced systems.
@@
Current approaches to reasoning about model performance degradation primarily focuses on isolated AI models, without considering the broader dynamic network environment. For example, drift detection [4] is commonly employed to identify performance degradation at the data level, detecting input changes and highlighting new or emerging data patterns. In addition, some heuristic methods [5] use historical data to infer potential causes of performance issues. While these methods can be somewhat effective, they face significant challenges in the context of 6G networks. First, the massive volume and high velocity of data within future networks can overwhelm traditional analytics, making real-time processing difficult. Second, the complexity of 6G network architectures and AI models introduces a degree of variability and unpredictability that these approaches cannot handle. Third, current methods often require extensive domain knowledge and manual intervention to interpret data and draw inferences, which limits their scalability and adaptability. Thus, there is an urgent need for more autonomous and intelligent reasoning mechanisms capable of overcoming these challenges in 6G.
@@
In this paper, we introduce an innovative LLM-powered Chain-of-Thought (CoT) method (LLM-CoT) for reasoning about AI performance degradation in 6G networks. The proposed LLM-CoT includes a “teacher” model, responsible for generating CoT rationales as training data, and a “student” model, which is a CoT model fine-tuned using this data. First, we employ a three-step zero-shot prompting to generate CoT rationales by a teacher LLM model, including network lecture generation, plan generation and rationale generation. The generated CoT rationales are utilized as teaching data that embody insights and reasoning aligned with network performance dynamics. Next, the CoT student model undergoes fine-tuning through a two-stage process. The first stage, “rationale generation teaching,” focuses on training the model to construct logical explanations for observed phenomena or data points. In the second stage, “answer inference teaching,” the model utilizes the previously generated rationales to infer answers to specific problems. During inference, the rationales produced in the first stage guide the model in making informed decisions based on test data, thereby enhancing its understanding and reasoning capabilities with respect to AI performance issues in 6G environments.
@@
We propose an innovative LLM-powered Chain-of-Thought (CoT) model designed for detailed reasoning tasks. Our method leverages zero-shot prompting of LLMs to generate rationale teaching data, which is then used to fine-tune a CoT student model, enabling it to perform sophisticated reasoning processes.
@@ II-A Reasoning AI Performance in Networks
Recent research has increasingly emphasized the deployment of AI models within communication networks [7, 8], highlighting the critical need for robust monitoring and reasoning frameworks. However, current efforts in reasoning AI performance primarily focus on the models in isolation, neglecting the dynamic and complex nature of the network environment in which these models operate. For example, Liu et al. [4] analyzed concept drift of input data and proposed an equal intensity k-means approach for its detection. Ahn et al. [5] proposed a heuristics-based model to reason about performance issues with multiple causes and effects.
While these approaches address drift detection and heuristic learning at the data and model design levels, they do not incorporate network analysis into their reasoning processes. This omission significantly lower the effectiveness of root cause analysis, particularly in the context of modern and next-generation communication networks.
@@ II-B LLMs for Intelligent Networks
As AI-native 6G networks evolve, integrating LLMs into network infrastructures is poised to become a crucial component, driving the next wave of autonomous network systems. The unique capabilities of LLMs, including their scalability and adaptability, make them indispensable for advancing network intelligence and efficiency, marking a new era in intelligent network evolution. In this paper, we propose an LLM-CoT model that significantly advances the AI reasoning for 6G. It leverages the depth and adaptability of LLMs to analyze AI performance and 6G networks, providing a promising solution to accurately identify the roots of performance degradation across network and model design factors.
@@ III-A2 Edge AI Models
For real-time 3D rendering tasks, we choose multi-Access Technologies (mATs) for data transmission. Specifically, the mATs encompass three wireless accesses including WiFi, 5G and LiFi, which can leverage the unique advantages of each access technology to ensure high throughput and low-latency transmission of volumetric data, facilitating distributed volumetric video capturing and real-time rendering. The mATs emphasize the importance of high reliability and connectivity required to enable next-generation digital services and experiences in 6G networks.
@@ III-C Problem Formulation for Reasoning
This paper addresses the issue of reasoning AI performance degradation in complex 6G network environments. The analysis focuses on an edge computing task, 3D rendering, as discussed in Section III-A1. The AI model used for this task is the 3D-GS model, described in detail in Section III-A2. The 6G network environment is characterized by its use of mATs-featured wireless connections, as outlined in Section III-B.
The reasoning framework employed is the Multimodal-CoT (MM-CoT) [15]. In this framework, the model, denoted as , processes multi-modal inputs: the language input and the vision input . The language input defines the reasoning questions. The question format is formulated as: “Question: Which of the following is most likely to cause degradation in 3D rendering performance?” and “Options: (A)-(G) mATs-based factors; (H) AI Model”. Furthermore, we include the context information about the mATs environment into the question as shown in Fig. 2. The vision input comprises image features extracted from multi-view images that are used in the real-time 3D rendering.
The output of the reasoning process is twofold: rationale generation and answer inference. The rationale provides the step-by-step problem analysis, i.e., chain-of-thought. The answer inference identifies the factors that lead to the AI performance degradation, which may include network-related or internal model issues.
@@ IV-B Teacher Model
Given the complexity of 6G networks and the extensive use of AI within the networks, relying on human-annotated CoT signals is impractical. Manual annotation is not only time-consuming but also demands high-level expertise from human annotators. To address these challenges, this paper adopts LLMs to generate CoT rationales via zero-shot prompting. Specifically, we employ LLM GPT-3.5 as the teacher model for generating teaching data. While GPT-3.5 possesses extensive knowledge across various fields, fully and accurately understanding AI operations within complex 6G network environments remains a significant challenge. Consequently, we adopt a zero-shot plan-and-solve prompting strategy [16] to generate PCoT. This approach innovatively designs a three-step zero-shot prompting for PCoT generation, tailored to 6G mATs and the edge AI model for 3D rendering.
@@ V-A Experimental Setup
Edge Task Simulation. To capture real-time images, we adopt four datasets for mATs-based image transmission: “Playroom” (900 images) and “DrJohnson” (1052 images) from the Deep Blending dataset [20], and “Truck” (1004 images) and “Train” (1204 images) from the Tank&Temples dataset [14]. We employ three access technologies for transmission—WiFi, 5G, and LiFi—with average throughputs of 800 Mbps, 400 Mbps, and 200 Mbps, respectively. Images are received from cameras and converted to bits, which are then processed using Quadrature Frequency Shift Keying (QFSK) for modulation and demodulation to achieve RGB image transmissions. These channels are randomly subject to Additive White Gaussian Noise (AWGN), which can distort the signal and impact the quality of transmitted images, thereby affecting the 3D rendering outcomes of the edge AI model 3D-GS.
@@
Baselines. First, we include API-based OpenAI LLMs (GPT-3.5 and GPT-4) for comparison, which are used zero-shot prompting for answer generation without fine-tuning. The prompting strategies of the two OpenAI LLMs follow the settings of [19, 23]. Next, two state-of-the-art (SoTA) fine-tuning based CoT models are used for comparison, including UnifiedQA-CoT [19] and Multimodal-CoT [15].
@@ VI Conclusion
In this study, we pioneered the investigation into the performance degradation of AI in 6G networks using LLMs. Unlike previous efforts that focused solely on AI model structures, our approach considered the AI model within a 6G-enabled mATs network environment, taking into account the effects of various access technologies for reasoning analytics. We selected 3D rendering as the edge AI task with mATs for data transmission and introduced the LLM-CoT method for in-depth reasoning. A teacher model was developed to generate teaching data, and a student model was subsequently fine-tuned with this data for improved rationale and answer prediction. Furthermore, we created a dataset, named AI-DR, specifically for reasoning about AI performance degradation in 6G contexts. Experimental results affirmed the validity of our dataset and the effectiveness of the LLM-CoT method.

### v2

@@ Abstract
6G, AI Performance Reasoning, Large Language Model, Chain-of-Thought, Multi-Access TechnologyI Introduction
As technology continues to evolve, network infrastructure is rapidly advancing towards the next-generation framework, notably 6G. Unlike its predecessors, 6G is expected to feature a deep and pervasive integration of Artificial Intelligence (AI) [1]. The native AI integration has the potential to significantly enhance the functionality and efficiency of 6G networks, offering powerful capabilities in terms of connectivity, reliability, and intelligent decision-making. However, the performance of AI models within these networks is critical. Any degradation in AI performance can significantly affect network operations and the services they support. Various factors can contribute to AI performance degradation, such as data quality issues, algorithmic biases, and sudden changes in network conditions. Moreover, the complex and dynamic nature of 6G can deteriorate these challenges, as AI models may struggle to adapt swiftly to the high-speed, voluminous data flows that characterize such advanced systems.
@@
Current approaches to reasoning about model performance degradation primarily focus on isolated AI models, without considering the broader dynamic network environment. For example, drift detection [4] is commonly employed to identify performance degradation at the data level, detecting input changes and highlighting new or emerging data patterns. In addition, some heuristic methods [5] use historical data to infer potential causes of performance issues. While these methods can be somewhat effective, they face significant challenges in the context of 6G networks. First, the massive volume and high velocity of data within future networks can overwhelm traditional analytics, making real-time processing difficult. Second, the complexity of 6G network architectures and AI models introduces a degree of variability and unpredictability that these approaches cannot handle. Third, current methods often require extensive domain knowledge and manual intervention to interpret data and draw inferences, which limits their scalability and adaptability. Thus, there is an urgent need for more autonomous and intelligent reasoning mechanisms capable of overcoming these challenges in 6G.
@@
In this paper, we introduce an innovative LLM-powered Chain-of-Thought (CoT) method (LLM-CoT) for reasoning about AI performance degradation in 6G networks. The proposed LLM-CoT includes a “teacher” model, responsible for generating CoT rationales as training data, and a “student” model, which is a CoT model fine-tuned using this data. First, we employ a three-step zero-shot prompting to generate CoT rationales by a teacher LLM model, including network lecture generation, plan generation and rationale generation. The generated CoT rationales are used as teaching data that embody insights and reasoning aligned with network performance dynamics. Next, the CoT student model undergoes fine-tuning through a two-stage process. The first stage, “rationale generation teaching,” focuses on training the model to construct logical explanations for observed phenomena or data points. In the second stage, “answer inference teaching,” the model employs the previously generated rationales to infer answers to specific problems. During inference, the rationales produced in the first stage guide the model in making informed decisions based on test data, thereby enhancing its understanding and reasoning capabilities with respect to AI performance issues in 6G environments.
@@
We propose an innovative LLM-powered CoT model designed for detailed reasoning tasks. Our method leverages zero-shot prompting of LLMs to generate rationale teaching data, which is then used to fine-tune a CoT student model, enabling it to perform sophisticated reasoning processes.
@@ II-A Reasoning AI Performance in Networks
Recent research has increasingly emphasized the deployment of AI models within communication networks [7, 8], highlighting the critical need for robust monitoring and reasoning frameworks. However, current efforts in reasoning AI performance primarily focus on the models in isolation, neglecting the dynamic and complex nature of the network environment in which these models operate. For example, Liu et al. [4] analyzed the concept drift of input data and proposed an equal intensity k-means approach for its detection. Ahn et al. [5] proposed a heuristics-based model to reason about performance issues with multiple causes and effects.
While these approaches address drift detection and heuristic learning at the data and model design levels, they do not incorporate network analysis into their reasoning processes. This omission significantly lowers the effectiveness of root cause analysis, particularly in the context of modern and next-generation communication networks.
@@ II-B LLMs for Intelligent Networks
As AI-native 6G networks evolve, integrating LLMs into network infrastructures is poised to become a crucial component, driving the next wave of autonomous network systems. The scalability and adaptability of LLMs make them essential for advancing network intelligence and efficiency, leading to a new era of intelligent network evolution. In this paper, we propose an LLM-CoT model that significantly advances the AI reasoning for 6G. It leverages the depth and adaptability of LLMs to analyze AI performance and 6G networks, providing a promising solution to accurately identify the roots of performance degradation across networks and model design factors.
@@ III-A2 Edge AI Models
For real-time 3D rendering tasks, we use multi-Access Technologies (mATs) for data transmission. Specifically, the mATs encompass three wireless accesses including WiFi, 5G and LiFi, which can leverage the unique advantages of each access technology to ensure high throughput and low-latency transmission of volumetric data. The mATs emphasize the importance of high reliability and connectivity required to enable next-generation digital services and experiences in 6G.
@@ III-C Problem Formulation for Reasoning
This paper addresses the issue of reasoning AI performance degradation in complex 6G network environments. The analysis focuses on an edge computing task, 3D rendering, as discussed in Section III-A1. The AI model used for this task is the 3D-GS model, as detailed in Section III-A2. The 6G network environment is characterized by its use of mATs-featured wireless connections, as outlined in Section III-B.
The reasoning framework employed is the Multimodal-CoT (MM-CoT) [15]. In this framework, the model denoted as , processes multi-modal inputs: the language input and the vision input . The language input defines the reasoning questions. The question format is formulated as: “Question: Which of the following is most likely to cause degradation in 3D rendering performance?” and “Options: (A)-(G) mATs-based factors; (H) AI Model”. Furthermore, we include the context information about the mATs environment into the question as shown in Fig. 2. The vision input comprises image features extracted from multi-view images that are used in the real-time 3D rendering.
The output of the reasoning process is twofold: rationale generation and answer inference. The rationale provides the step-by-step problem analysis, i.e., chain-of-thought. The answer inference identifies the factors that lead to AI performance degradation, which may include network-related or internal model issues.
@@ IV-B Teacher Model
Given the complexity of 6G networks and their extensive use of AI, relying on human-annotated CoT signals is impractical. Manual annotation is not only time-consuming but also demands high-level expertise from human annotators. To address these challenges, this paper adopts LLMs to generate CoT rationales via zero-shot prompting. Specifically, we employ LLM GPT-3.5 as the teacher model for generating teaching data. While GPT-3.5 possesses extensive knowledge across various fields, fully and accurately understanding AI operations within complex 6G network environments remains a significant challenge. Consequently, we adopt a zero-shot plan-and-solve prompting strategy [16] to generate PCoT. This approach innovatively designs a three-step zero-shot prompting for PCoT generation, tailored to 6G mATs and the edge AI model for 3D rendering.
@@ V-A Experimental Setup
Edge Task Simulation. To capture real-time images, we adopt four datasets for mATs-based image transmission: “Playroom” (900 images) and “DrJohnson” (1052 images) from the Deep Blending dataset [20], and “Truck” (1004 images) and “Train” (1204 images) from the Tank&Temples dataset [14]. We employ three access technologies for transmission-WiFi, 5G, and LiFi-with average throughputs of 800 Mbps, 400 Mbps, and 200 Mbps, respectively. Images are received from cameras and converted to bits, which are then processed using Quadrature Frequency Shift Keying for modulation and demodulation to achieve RGB image transmissions. These channels are randomly subject to Additive White Gaussian Noise, which can distort the signal and impact the quality of transmitted images, thereby affecting the 3D rendering outcomes of the edge AI model 3D-GS.
@@
Baselines. First, we include API-based OpenAI LLMs (GPT-3.5 and GPT-4) for comparison, which use zero-shot prompting for answer generation without fine-tuning. The prompting strategies of the two OpenAI LLMs follow the settings of [19, 23]. Next, two state-of-the-art (SoTA) fine-tuning based CoT models are used for comparison, including UnifiedQA-CoT [19] and Multimodal-CoT [15].
@@ VI Conclusion
In this study, we pioneered the investigation into the performance degradation of AI in 6G networks using LLMs. Unlike previous efforts that focused solely on AI model structures, our approach considered the AI model within a 6G-enabled mATs network environment, taking into account the effects of various access technologies for reasoning analytics. We selected 3D rendering as the edge AI task with mATs for data transmission and introduced the LLM-CoT method for in-depth reasoning. A teacher model was developed to generate teaching data, and a student model was subsequently fine-tuned with this data for improved rationale and answer prediction. Furthermore, we created a dataset, named AI-DR, specifically for reasoning about AI performance degradation in 6G contexts. Experimental results confirmed the validity of our dataset and the effectiveness of the LLM-CoT method.